OpenStudy (anonymous):

Decide whether each equation has one solution, no solutions, or infinitely many solutions. 4(x + 3) + 2x = x – 8

OpenStudy (anonymous):

did u try solving it yet?

OpenStudy (anonymous):

I need help on how to solve it

OpenStudy (anonymous):

\[a(x+b)=ax+ab\]\[ax+bx=(a+b)x\]

OpenStudy (anonymous):

\[ax=bx+c \rightarrow ax-bx=c\]

OpenStudy (anonymous):

\[ax=c \rightarrow x=\frac{ c }{ a }\]

OpenStudy (anonymous):

I don't get it.

OpenStudy (anonymous):

where a, b, c are constant real numbers

OpenStudy (anonymous):

follow these steps to solve the equation. Memorize them

OpenStudy (anonymous):

A 4 B 3 C 2 4(x+3)=4x+43 4x+3x=(4+3)x 4x=3x+2->4x-3x=2 4x=2->x=2/4